King's Road recently acquired all of Oxford Corporation's stock and is now consolidating the financial data of this new subsidiary. King's Road paid a total of $1,020,000 for Oxford, which has the following accounts:

Fair Value Tax Basis
Accounts receivable $ 210,000 $ 210,000
Inventory 192,000 192,000
Land 175,000 175,000
Buildings 187,000 142,000
Equipment 222,750 154,000
Liabilities (223,000 ) (223,000 )
  1. What amount of deferred tax liability arises in the acquisition?

  2. What amounts will be used to consolidate Oxford with King's Road at the date of acquisition?

  3. On a consolidated balance sheet prepared immediately after this takeover, how much goodwill should King's Road recognize? Assume a 40 percent effective tax rate.
Answer #1

a.

Deferred tax liability arises from temporary differences. DTL is symbol of taxes payable in the future and the taxes that are currently due. Temporary differences are capable of reversal in the future.

Differences in

Building = 187,000 – 142,000 = 45,000

Equipment = 222,750 – 154,000 = 68,750

Net assets difference = 113,750

Deferred tax liability = difference between fair value and tax value as above X 40% = 113,750*40% = 45,500

b.Amounts to be used to consolidate at time of acquisition

Accounts receivable

210,000

Inventory

192,000

Land

175,000

Building

187,000

Equipment

222,750

Total assets

986,750

Less: liabilities

223,000

Net assets

763,750

c.

Computation of goodwill to be recognized on consolidated balance sheet.

Particulars

Amount ($)

Amount paid to acquire company

1020,000

Value of assets taken over                           986,750

Less: liability                                                   223,000

Less: deferred tax liability                             45,500

Net assets taken over

718,250

Goodwill

301,750

Hence, goodwill to be recognized would be $ 301,750

kindly upvote